OpenStudy (anonymous):

express as a quotient x^-4+x^6

OpenStudy (anonymous):

\[\frac{1}{x^4}+x^6\] is a start

OpenStudy (anonymous):

write over one denominator via \[\frac{1}{x^4}+\frac{x^6\times x^4}{x^4}\]\[=\frac{1+v^{10}}{v^4}\]

OpenStudy (kropot72):

@satellite73 means the answer is\[\frac{1+x ^{10}}{x ^{4}}\]

OpenStudy (anonymous):

yeah i guess so not sure where the \(v\) came from

OpenStudy (kropot72):

The v came from typo-world surely :)